A company reports the following beginning inventory and two purchases for the month of January. On January 26, the company sells 400 units. Ending inventory at January 31 totals 150 units. Units Unit Cost Beginning inventory on January 1 360 $ 3.50 Purchase on January 9 80 3.70 Purchase on January 25 110 3.80 Required: Assume the perpetual inventory system is used. Determine the costs assigned to ending inventory when costs are assigned based on the weighted average method. (Round your per unit costs to 2 decimal places.)
Answer #1

Calculate ending inventory and cost of goods sold

Unit Unit Cost Total Cost
Beginning inventory 360 3.50 1260
Purchase 9 80 3.70 296
Purchase 25 110 3.80 418
Total 550 1974

Cost of goods sold = (1974/550) = 3.59 *400 = $1436

Ending inventory = 3.59*150 = 538.50